# Customizing Visual Emotion Evaluation for MLLMs: An Open-vocabulary, Multifaceted, and Scalable Approach

### v1

@@ 1 Introduction
When applied to MLLMs, these methods reveal four primary limitations. Firstly, their adoption of fixed ground-truth answers for open-ended questions imposes structural constraints that exclude other plausible responses. Emotion perception is inherently subjective (Zhao et al., 2016), as the same image may evoke divergent reactions across individuals, and emotional states permit varied interpretations. As demonstrated in Figure˜1, responses generated by GPT4o (Hurst et al., 2024) that seem reasonable to humans are judged as inaccurate under rigid evaluation metrics. Secondly, they are mostly constructed upon emotion theories with limited emotional taxonomies. Popular emotion classification and interpretation benchmarks, such as FI (You et al., 2016) and Artemis (Achlioptas et al., 2021), comprise only eight emotion categories. Such taxonomic granularity fails to capture fine-grained affective variations between images. Thirdly, they focus solely on intrinsic image attributes while overlooking critical contextual dimensions. As recognized in established psychological literature, emotion perception can be influenced by extravisual factors (Barrett et al., 2011), including the scene context in which the image is set (Wieser & Brosch, 2012), as well as the viewer’s identity and personality (Hamann & Canli, 2004). Fourthly, they predominantly rely on majority voting mechanisms to ensure label reliability in crowdsourced annotations (Li et al., 2017), which is labor-intensive, particularly for fine-grained annotation tasks. EMOTIC (Kosti et al., 2017), for instance, requires coordination with 23,788 annotators. This operational burden severely constrains dataset scalability in magnitude and generalization capacity across image domains.
@@
Leveraging INSETS, we automatically construct INSETS-462k, a large-scale annotated ESJ corpus. Building on it, we curate MVEI benchmark (Multifaceted evaluation of Visual Emotion Intelligence) through careful human refinement. MVEI comprises 3,086 unique image–statement pairs designed to enable comprehensive evaluation of MLLMs. Grounded in established theories of affective cognition, it covers four complementary dimensions: sentiment polarity (Russell, 1980), emotion interpretation (Ekman & Friesen, 1971), scene context (Barrett et al., 2011), and perception subjectivity (Hamann & Canli, 2004). Systematic evaluation reveals that recent MLLMs exhibit considerable proficiency but still lag behind humans, particularly in discerning emotional polarity and interpreting perception subjectivity. Further explorations indicate that the former can likely be improved through targeted adaptation, whereas the latter is more tied to the models’ inherent properties, highlighting potential directions for future research. In summary, the contributions of this paper are threefold:
@@ 2.1 AICA Benchmarks
Psychological researchers conceptualize emotion representation through two principal frameworks: the Categorical Emotion Space (CES), which discretizes affective states into predefined taxonomies, and the Dimensional Emotion Space (DES), which maps emotions onto continuous coordinations. For simplicity and better interpretability, most benchmarks adopt emotion classification evaluations based on discrete CES emotion taxonomies. This category encompasses both early small-scale benchmarks, such as IAPSa (Mikels et al., 2005) and Abstract (Machajdik & Hanbury, 2010), as well as later larger-scale benchmarks like FI (You et al., 2016) and WebEmo (Panda et al., 2018). Over time, benchmarks with enriched metadata have also been developed. Notable examples include EMOTIC (Kosti et al., 2017), which integrates multiple emotion categories, VAD values (Schlosberg, 1954), and human-related bounding boxes, and EmoSet (Yang et al., 2023a), which employs describable emotion attributes that cover different levels of visual information.
Some other benchmarks adopt emotion interpretation evaluations by extending CES-based taxonomies with additional emotional explanations, such as Artemis (Achlioptas et al., 2021) and Affection (Achlioptas et al., 2023). EIBench (Lin et al., 2025) diverges slightly, shifting focus on identifying visual emotional triggers. Based on these benchmarks, numerous expert models (Xu et al., 2022; Jia & Yang, 2022; Feng et al., 2023) have been developed, demonstrating strong performance under the fine-tuning and testing paradigm. In contrast, MLLMs are commonly pre-trained on web-scale data, without explicitly aligning with benchmark-specific knowledge. This discrepancy introduces multiple constraints when applying conventional benchmarks to MLLMs, necessitating customized visual emotion evaluation approaches that account for their generalized knowledge structures.
@@ 2.2 Evaluation of MLLMs
Recent years have witnessed growing academic and industrial interest in MLLMs. Unlike specialized models, MLLMs demonstrate versatile competence across diverse tasks (Li & Lu, 2024), fueling expectations for their trajectory toward Artificial General Intelligence (Maruyama, 2020). To evaluate MLLMs, various benchmarks have been established, covering perception (Li et al., 2023b; Liu et al., 2023), reasoning (Nie et al., 2024; Zhang et al., 2019), ethics (Qian et al., 2024b; Guan et al., 2024), and specialized domains (Chen et al., 2024a; Qian et al., 2024a). Yet emotional intelligence remains conspicuously underexplored, particularly in the visual modality. In existing efforts, MM-BigBench (Yang et al., 2023b) simply aggregates mainstream image-text benchmarks; FABA-Bench (Li et al., 2024) focuses primarily on facial expressions and actions; and EmoBench-M (Hu et al., 2025) has not escaped from coarse-grained emotion taxonomy. To fill this gap, we propose the ESJ task and the MVEI benchmark, aiming to customize and advance visual emotion evaluation of MLLMs.
@@ 3 Emotion Statement Judgement
Collectively, these dimensions establish a holistic visual emotion evaluation framework for MLLMs. They cover both intrinsic image attributes emphasized in existing benchmarks and underexplored contextual factors critical for human emotional perception (Stemmler & Wacker, 2010).
@@ 4.1 Open-vocabulary Emotion Tagging
At this stage, INSETS aims to assign open-vocabulary emotion labels for images, laying a solid foundation for constructing meaningful emotion-centric statements, with its procedure depicted in Figure˜2. According to Cheng et al. (2024), MLLMs demonstrate promising capabilities in generating emotional descriptions from visual content and extracting underlying emotions from these descriptions. However, challenges such as hallucinations (Bai et al., 2024), trustworthiness issues (Liu et al., 2024b), and inherent limitations in emotional perception can lead to inaccuracies in the extracted emotions. To enhance reliability, we devise an ensemble-based majority voting mechanism, aggregating outputs from multiple MLLMs to cross-validate and refine emotion label assignments.
@@ 4.3 Construction of INSETS-462k and MVEI
Table 3: Statistics of the human refinement process.
@@ 5.3 Results and Findings
Qwen2.5-VL (8.3B) #Shots Accuracy (%)
@@
In-Context Learning: Random Retrieval 2 66.3 (+3.1) 81.6 (+0.1) 84.8 (+0.9) 66.5 (+0.2) 76.6 (+0.7)
@@
LoRA Fine-Tuning (Hu et al., 2022) - 78.6 (+15.4) 84.7 (+3.2) 86.3 (+2.4) 70.3 (+4.0) 82.0 (+5.4)
Full Parameter Fine-Tuning (Freeze Vision) - 84.3 (+21.1) 84.8 (+3.3) 87.0 (+3.1) 71.1 (+4.8) 83.3 (+6.7)
GRPO (Shao et al., 2024) - 83.2 (+20.0) 82.5 (+1.0) 86.5 (+2.6) 71.1 (+4.8) 82.0 (+5.4)
@@
As shown in Table˜6, all applied techniques consistently improve MLLM performance, demonstrating the benefits of both in-context learning and task-specific fine-tuning. The most pronounced gains occur in sentiment polarity, indicating that MLLMs possess the capability to capture overall emotional tone. Their previous deficiency is likely due to confusion between positive, negative, and mixed categories, and it can be effectively alleviated through few-shot demonstrations or lightweight fine-tuning. By contrast, perception subjectivity shows only modest improvement and remains the weakest dimension, reflecting a more fundamental challenge that may require subjectivity-oriented training objectives or specialized datasets. While targeted adaptation on INSETS-462k provides clear benefits, we view this work primarily as a foundational benchmark for advancing emotional intelligence in more general-purpose MLLMs. Rather than treating ESJ as a direct optimization target, we advocate its use as an evaluation metric and feedback signal to guide broader model development.
@@ 7 Ethics Statement
This study evaluates MLLMs in visual emotion perception through a new task, pipeline, and benchmark. While these contributions aim to advance research, several ethical considerations merit attention. First, since MLLMs are typically pre-trained on large-scale, publicly available but often uncurated datasets, they may inadvertently inherit biases or sensitive content that could also be reflected in the automatically constructed and unrefined INSETS-462K corpus. We therefore recommend that researchers employ the corpus with appropriate caution and awareness of potential dataset biases. Second, the next token prediction nature of MLLMs can occasionally yield inaccurate or misleading outputs, emphasizing the need to avoid uncritical reliance on such predictions in high-stakes or sensitive contexts.
All data used in this work are drawn from publicly available benchmarks, and no private information was collected or disclosed. By acknowledging these considerations, we aim to promote responsible use of our data, mitigate potential risks, and support its positive impact on future research.
@@ Appendix D Details of Parrott’s Hierarchical Model
Appendix E Visualization of MVEI
More samples from MVEI are visualized in Figure˜5, Figure˜6, Figure˜7, Figure˜8, Figure˜9, Figure˜10, Figure˜11, Figure˜12.
Figure 5: Correct sentiment polarity statements.
Figure 6: Incorrect sentiment polarity statements.
Figure 7: Correct emotion interpretation statements.
Figure 8: Incorrect emotion interpretation statements.
Figure 9: Correct scene context statements.
Figure 10: Incorrect scene context statements.
Figure 11: Correct perception subjectivity statements.
Figure 12: Incorrect perception subjectivity statements.
Appendix F Links of MLLMs

### v2

@@ 1 Introduction
When applied to MLLMs, these methods reveal four primary limitations. Firstly, their adoption of fixed ground-truth answers for open-ended questions imposes structural constraints that exclude other plausible responses. Emotion perception is inherently subjective (Zhao et al., 2016), as the same image may evoke divergent reactions across individuals, and emotional states permit varied interpretations. As demonstrated in Figure˜1, responses generated by GPT4o (Hurst et al., 2024) that seem reasonable to humans are judged as inaccurate under rigid evaluation metrics. Secondly, they are mostly constructed upon emotion theories with limited emotional taxonomies. Popular emotion classification and interpretation benchmarks, such as FI (You et al., 2016) and Artemis (Achlioptas et al., 2021), comprise only eight emotion categories. Such taxonomic granularity fails to capture fine-grained affective variations between images. Thirdly, they focus solely on intrinsic image attributes while overlooking critical contextual dimensions. As recognized in established psychological literature, emotion perception can be influenced by extravisual factors (Barrett et al., 2011), including the scene context in which the image is set (Wieser and Brosch, 2012), as well as the viewer’s identity and personality (Hamann and Canli, 2004). Fourthly, they predominantly rely on majority voting mechanisms to ensure label reliability in crowdsourced annotations (Li et al., 2017), which is labor-intensive, particularly for fine-grained annotation tasks. EMOTIC (Kosti et al., 2017), for instance, requires coordination with 23,788 annotators. This operational burden severely constrains dataset scalability in magnitude and generalization capacity across image domains.
@@
Leveraging INSETS, we automatically construct INSETS-462k, a large-scale annotated ESJ corpus. Building on it, we curate MVEI benchmark (Multifaceted evaluation of Visual Emotion Intelligence) through careful human refinement. MVEI comprises 3,086 unique image–statement pairs designed to enable comprehensive evaluation of MLLMs. Grounded in established theories of affective cognition, it covers four complementary dimensions: sentiment polarity (Russell, 1980), emotion interpretation (Ekman and Friesen, 1971), scene context (Barrett et al., 2011), and perception subjectivity (Hamann and Canli, 2004). Systematic evaluation reveals that recent MLLMs exhibit considerable proficiency but still lag behind humans, particularly in discerning emotional polarity and interpreting perception subjectivity. Further explorations indicate that the former can likely be improved through targeted adaptation, whereas the latter is more tied to the models’ inherent properties, highlighting potential directions for future research. In summary, the contributions of this paper are threefold:
@@ 2.1 AICA Benchmarks
Psychological researchers conceptualize emotion representation through two principal frameworks: the Categorical Emotion Space (CES), which discretizes affective states into predefined taxonomies, and the Dimensional Emotion Space (DES), which maps emotions onto continuous coordinations. For simplicity and better interpretability, most benchmarks adopt emotion classification evaluations based on discrete CES emotion taxonomies. This category encompasses both early small-scale benchmarks, such as IAPSa (Mikels et al., 2005) and Abstract (Machajdik and Hanbury, 2010), as well as later larger-scale benchmarks like FI (You et al., 2016) and WebEmo (Panda et al., 2018). Over time, benchmarks with enriched metadata have also been developed. Notable examples include EMOTIC (Kosti et al., 2017), which integrates multiple emotion categories, VAD values (Schlosberg, 1954), and human-related bounding boxes, and EmoSet (Yang et al., 2023a), which employs describable emotion attributes that cover different levels of visual information.
Some other benchmarks adopt emotion interpretation evaluations by extending CES-based taxonomies with additional emotional explanations, such as Artemis (Achlioptas et al., 2021) and Affection (Achlioptas et al., 2023). EIBench (Lin et al., 2025) diverges slightly, shifting focus on identifying visual emotional triggers. Based on these benchmarks, numerous expert models (Jia and Yang, 2022; Feng et al., 2023; Wu et al., 2024) have been developed, demonstrating strong performance under the fine-tuning and testing paradigm. In contrast, MLLMs are commonly pre-trained on web-scale data, without explicitly aligning with benchmark-specific knowledge. This discrepancy introduces multiple constraints when applying conventional benchmarks to MLLMs, necessitating customized visual emotion evaluation approaches that account for their generalized knowledge structures.
@@ 2.2 Evaluation of MLLMs
Recent years have witnessed growing academic and industrial interest in MLLMs. Unlike specialized models, MLLMs demonstrate versatile competence across diverse tasks (Li and Lu, 2024), fueling expectations for their trajectory toward Artificial General Intelligence (Maruyama, 2020). To evaluate MLLMs, various benchmarks have been established, covering perception (Li et al., 2023b; Liu et al., 2023), reasoning (Nie et al., 2024; Zhang et al., 2019), ethics (Qian et al., 2024b; Guan et al., 2024), and specialized domains (Chen et al., 2024a; Qian et al., 2024a). Yet emotional intelligence remains conspicuously underexplored, particularly in the visual modality. In existing efforts, MM-BigBench (Yang et al., 2023b) simply aggregates mainstream image-text benchmarks; FABA-Bench (Li et al., 2024) focuses primarily on facial expressions and actions; EmoBench-M (Hu et al., 2025) and EEmo-Bench (Gao et al., 2025), while largely extending task coverage, still insufficiently handle the ambiguity inherent in open-ended questions. To fill this gap, we propose the ESJ task and the MVEI benchmark, aiming to customize and advance visual emotion evaluation of MLLMs.
@@ 3 Emotion Statement Judgement
Collectively, these dimensions establish a holistic visual emotion evaluation framework for MLLMs. They cover both intrinsic image attributes emphasized in existing benchmarks and underexplored contextual factors critical for human emotional perception (Stemmler and Wacker, 2010).
@@ 4.1 Open-vocabulary Emotion Tagging
At this stage, INSETS aims to assign open-vocabulary emotion labels for images, laying a solid foundation for constructing meaningful emotion-centric statements, with its procedure depicted in Figure˜2. According to (Cheng et al., 2024), MLLMs demonstrate promising capabilities in generating emotional descriptions from visual content and extracting underlying emotions from these descriptions. However, challenges such as hallucinations (Bai et al., 2024), trustworthiness issues (Liu et al., 2024b), and inherent limitations in emotional perception can lead to inaccuracies in the extracted emotions. To enhance reliability, we devise an ensemble-based majority voting mechanism, aggregating outputs from multiple MLLMs to cross-validate and refine emotion label assignments.
@@ 4.3 Construction of INSETS-462k and MVEI
Table 3: Statistics of the human refinement process. Kappa represents Fleiss’ Kappa.
@@ 5.3 Results and Findings
Qwen2.5-VL (8.3B) #Shot Accuracy (%)
@@
Chain-of-Thought Reasoning - 67.4 (+4.2) 81.5 (+0.0) 84.6 (+0.7) 67.0 (+0.7) 76.6 (+0.8)
In-Context Learning: Random Retrieval 2 66.3 (+3.1) 81.6 (+0.1) 84.8 (+0.9) 66.5 (+0.2) 76.9 (+1.0)
@@
LoRA Fine-Tuning (Hu et al., 2022) - 78.6 (+15.4) 84.7 (+3.2) 86.3 (+2.4) 70.3 (+4.0) 80.7 (+4.8)
Full Parameter Fine-Tuning (Freeze Vision) - 84.3 (+21.1) 84.8 (+3.3) 87.0 (+3.1) 71.1 (+4.8) 81.9 (+6.0)
GRPO (Shao et al., 2024) - 83.2 (+20.0) 82.5 (+1.0) 86.5 (+2.6) 71.1 (+4.8) 80.7 (+4.8)
@@
As shown in Table˜6, all applied techniques consistently improve MLLM performance, demonstrating the benefits of both in-context learning and task-specific fine-tuning. The most pronounced gains occur in sentiment polarity, indicating that MLLMs possess the capability to capture overall emotional tone. Their previous deficiency is likely due to confusion between positive, negative, and mixed categories, and it can be effectively alleviated through few-shot demonstrations or lightweight fine-tuning. By contrast, perception subjectivity shows only modest improvement and remains the weakest dimension, reflecting a more fundamental challenge that may require subjectivity-oriented pre-training objectives or specialized datasets. While targeted adaptation on INSETS-462k provides clear benefits, we view this work primarily as a foundational benchmark for advancing emotional intelligence in more general-purpose MLLMs. Rather than treating ESJ as a direct optimization target, we advocate its use as an evaluation metric and feedback signal to guide broader model development.
@@ 7 Ethics Statement
This study evaluates MLLMs in visual emotion comprehension through a new task, pipeline, and benchmark. While we aim to advance research, several ethical considerations merit attention.
First, the dataset exhibits certain distributional imbalances in the emotion label of images. The images used in this work originate from EmoSet, which ultimately traces back to user-generated posts on social media. Such posts naturally reflect platform-specific emotional biases, most notably, positive content typically appears more frequently than negative content (Niu et al., 2016). This characteristic carries over to our MVEI benchmark, where, as shown in Figure˜4, positive-polarity images account for 65.2% of the data, compared to 34.8% for negative ones. While this skew partially mirrors real-world content distributions, it may influence model behavior or downstream analysis if not interpreted carefully. We therefore encourage users to remain aware of these imbalances to avoid biased or misleading conclusions.
Second, perception subjectivity statements may contain latent demographic biases arising from the automatically generated characters. Although extreme or inappropriate cases are filtered through human refinement, demographic attributes, such as age, gender, or cultural background, may still be unevenly reflected or stereotypically implied by the MLLMs. And since the characters are produced without explicit structural control, systematically quantifying their demographic statistics remains challenging. This limitation introduces the risk of subtle demographic skew being propagated or reinforced through the benchmark. Empirical quantification of these demographic patterns would enable finer-grained evaluation and customization, and we regard this as an important direction for future development.
Third, MLLM-generated data may still exhibit cultural-perspective and aesthetic-perception biases. Prior work has shown that cultural tendencies embedded in training corpora, such as language distribution and region-specific viewpoints, can be amplified during LLM inference, and even multilingual models often fail to equitably represent diverse cultural values (Tao et al., 2024). In addition, LLMs have been found to exhibit aesthetic preferences that may implicitly reinforce stereotypical standards (Kotek et al., 2023). Such tendencies could be potentially carried over into the constructed emotional labels and statements. Although the bias of any specific model can be alleviated through the proposed ensemble-based majority voting mechanism, these biases cannot be fully eliminated. Users should therefore interpret culturally sensitive results with caution and avoid overgeneralizing findings.
Fourth, annotation-level biases due to cultural differences or personal experiences may persist. Emotion perception is inherently subjective, and these differences can shape annotation outcomes. While ESJ task formulation targets specifically for such issues, it can hardly guarantee the complete elimination of these biases. Therefore, these concerns also warrant caution from users.
Although addressing these ethical concerns falls beyond the immediate scope of this study, we document them here to maintain transparency. We hope this clarifies the limitations of the benchmark and supports future efforts toward mitigation. All data used in this work are drawn from publicly available benchmarks, and no private information was collected or disclosed. By acknowledging these considerations, we aim to promote responsible use of our data, mitigate potential risks, and support its positive impact on future research.
8 Reproducibility Statement
To ensure reproducibility, the manuscript provides comprehensive documentation of the INSETS implementation, the human refinement process, and detailed statistics of both the INSETS-462K corpus and the MVEI benchmark. We release code and data on: https://github.com/wdqqdw/MVEI.
@@ Appendix D Details of Parrott’s Hierarchical Model
Appendix E Formalization of the Majority-Voting Mechanism
In this section, we provide a formalized definition of the majority-voting mechanism for clarification and transparency. Let an image be processed by $n$ MLLMs. The set of open-vocabulary emotion labels generated by the $i$-th model is denoted by $L_{i}=\{e_{i,1},e_{i,2},\dots,e_{i,m}\}$, where $i\in\{1,2,\dots,n\}$.
Let $\mathcal{P}$ be the mapping from an open-vocabulary emotion to the secondary emotion in POM. The ensemble-based majority-voting procedure is defined as follows.
  1. 1.
    Secondary-Emotion Quota. For a secondary emotion $\bar{e}$, define its quota $Q(\bar{e})$ as:
    $Q(\bar{e})=\max\left(\left\lfloor\sum_{i=1}^{n}\mathbb{I}[\bar{e}\in\{\mathcal{P}(e)\mid e\in L_{i}\}]-\frac{n}{2}+1\right\rfloor,0\right)$
    where $\mathbb{I}[\cdot]$ is the indicator function. This quantity is strictly positive only when $\bar{e}$ is supported by a majority of the models.
  2. 2.
    Candidate Pool Formation. The candidate pool of open-vocabulary labels for the secondary emotion $\bar{e}$ is defined as:
    $S(\bar{e})=\bigl\{\,e_{i,j}\,\bigm|\,i\in\{1,2,\dots,n\};\;j\in\{1,2,\dots,m\};\;\mathcal{P}(e_{i,j})=\bar{e}\,\bigr\}.$
  3. 3.
    Consensus Selection. Let $\mathrm{freq}(e)$ denotes the frequency of an open-vocabulary label $e$ in $S(\bar{e})$: $\mathrm{freq}(e)=|\{e_{i,j}\in S(\bar{e})|e_{i,j}=e\}|$. The consensus labels for $\bar{e}$ are defined as the top-$Q(\bar{e})$ unique labels in $S(\bar{e})$ ranked by $\mathrm{freq}(\cdot)$, where ties are resolved uniformly at random. The final consensus label set for the image is obtained by taking the union over all secondary emotions:
    $L_{\mathrm{cons}}=\bigcup_{\bar{e}}\mathrm{Top}_{Q(\bar{e})}\!\left(S(\bar{e})\right).$
Appendix F Visualization of Ambiguous Samples
Figure 5: Samples that are deemed ambiguous during the human refinement process.
Figure˜5 presents three representative samples that are identified as ambiguous during the manual refinement process and thus excluded from the MVEI benchmark. In the left case, the contextual description is considered an inappropriate or overextended inference from the visual content. In the middle case, the emotional interpretation is overly brief, with cues that were too general or vague to support a reliable emotional conclusion. In the right case, the characterization lacked sufficient specificity, where two distinct emotions remain equally plausible.
We attribute these ambiguities to two principal factors: the inherent limitations of MLLMs in visual perception, which manifest as inaccurate descriptions or superficial analyses, and the intrinsic challenges of emotion-related data construction, where contextual subjectivity can render multiple interpretations valid. These observations collectively highlight the critical necessity of human refinement in data construction pipelines involving MLLMs.
Appendix G Agreement between Assigned Labels and EmoSet Labels
To further evaluate the automatically assigned open-vocabulary labels, we perform a cross-validation based on the original EmoSet labels. Since EmoSet is built upon the Mikels model (Mikels et al., 2005), which contains eight emotion categories (amusement, awe, contentment, excitement, anger, disgust, fear, and sadness), it is not naturally aligned with our adopted Parrott hierarchical model.
However, we note that except for awe, the remaining seven Mikels categories are distributed across different levels of Parrott’s model. To enable comparison, we map them to the primary-level emotions in Parrott’s model: amusement → joy, contentment → joy, excitement → joy, anger → anger, disgust → anger, fear → fear, sadness → sadness. Finally, using the open-vocabulary attachment obtained from our constructed POM, we map awe → surprise. Based on this mapping, our analysis shows that the automatically assigned open-vocabulary labels of 97.3% of the 17,716 images in INSETS-462k overlap with their original EmoSet labels at the primary level of the Parrott model. This high level of consistency provides strong evidence for the validity of our automated annotation pipeline.
Appendix H Additional Results of MLLMs Adaptation Techniques
Table˜9 reports results of more comprehensive MLLM adaptation techniques on MVEI, building upon those in Table˜6. The 459k samples used for supervised fine-tuning correspond to the full INSETS-462k dataset with MVEI excluded, and the 50k samples used for reinforcement learning are a sampled subset of the former. For inference and GRPO, the CoT prompt are constructed by appending “think step-by-step” to the original prompt.
Table 9: Comprehensive Evaluation of MLLMs adaptation techniques on MVEI.
Qwen2.5-VL (8.3B) #Shot #Sample Accuracy (%)
Sentiment Emotion Scene Perception Total
Polarity Interpretation Context Subjectivity
Direct Inference - - 63.2 81.5 83.9 66.3 75.9
Chain-of-Thought (CoT) Reasoning - - 67.4 81.5 84.6 67.0 76.6
In-Context Learning: Random Retrieval 2 - 66.3 81.6 84.8 66.5 76.6
In-Context Learning: Random Retrieval 4 - 68.8 81.7 85.0 66.7 77.1
In-Context Learning: Random Retrieval 8 - 70.1 81.7 84.9 67.0 77.3
LoRA Fine-Tuning (Hu et al., 2022) - 10k 78.6 84.7 86.3 70.3 80.7
LoRA Fine-Tuning (Hu et al., 2022) - 459k 82.2 86.0 86.9 71.9 82.2
Full Parameter Fine-Tuning (Freeze Vision) - 10k 84.3 84.8 87.0 71.1 81.9
Full Parameter Fine-Tuning (Freeze Vision) - 459k 85.6 86.5 87.6 73.3 83.4
GRPO (Shao et al., 2024) (without CoT) - 10k 83.2 82.5 86.5 71.1 80.7
GRPO (Shao et al., 2024) (without CoT) - 50k 84.0 82.7 86.3 71.4 80.9
GRPO (Shao et al., 2024) (with CoT) - 10k 86.2 82.9 86.6 72.3 81.6
GRPO (Shao et al., 2024) (with CoT) - 50k 86.8 83.0 87.2 72.7 82.0
Appendix I Links of MLLMs